OpenStudy (anonymous):

Which equation is equivalent to –k + 0.03 + 1.01k = –2.45 – 1.81k?

OpenStudy (anonymous):

–100k + 3 + 101k = –245 – 181k –100k + 300 + 101k = –245 – 181k –k + 3 + 101k = –245 – 181k –k + 300 + 101k = –245 – 181k

OpenStudy (anonymous):

i got something totally different i got 0.01k+0.03=-2.45-1.81k

OpenStudy (anonymous):

this website should make it easier for you https://mathway.com/

OpenStudy (anonymous):

it honestly did not

OpenStudy (anonymous):

thanks though

OpenStudy (anonymous):

@159507nb

OpenStudy (anonymous):

@TheSmartOne

OpenStudy (dan815):

multiply left and right side by 100 what happens?

OpenStudy (anonymous):

u would get –100k + 3 + 101k = –245 – 181k

OpenStudy (anonymous):

so A